# Domain Adaptation from Generated Multi-Weather Images for Unsupervised Maritime Object Classification

### v2

@@
Abstract.
The classification and recognition of maritime objects are crucial for enhancing maritime safety, monitoring, and intelligent sea environment prediction. However, existing unsupervised methods for maritime object classification often struggle with the long-tail data distributions in both object categories and weather conditions. In this paper, we construct a dataset named AIMO produced by large-scale generative models with diverse weather conditions and balanced object categories, and collect a dataset named RMO with real-world images where long-tail issue exists. We propose a novel unsupervised domain adaptation approach that leverages AIMO (source domain) to address the problem of limited labeled data, unbalanced distribution and domain shift in RMO (target domain). Specifically,we enhance the generalization of source features with the Vision-Language Models (i.e., CLIP) by Self-Knowledge Distillation. And we propose a difficulty score for curriculum learning to optimize training process by gradually introducing the generated data. Experimental results shows that the proposed method significantly improves the classification accuracy, particularly for samples within rare object categories and weather conditions. Datasets and codes will be publicly available.
Domain Adaptation, Maritime Object Classification
copyright: acmlicensedjournalyear: 2025doi: XXXXXXX.XXXXXXXconference: The 33rd ACM International Conference on Multimedia; June 03–05, 2025; Woodstock, NYisbn: 978-1-4503-XXXX-X/2018/06ccs: Computing methodologies Object recognition
1. Introduction
@@
Unsupervised maritime object classification presents significant challenges in handling the high similarity between ships and complex maritime environments. Large-scale annotated datasets are not common because of the difficulty in obtaining maritime objects and the high cost of manual labeling (Wang et al., 2020). Salem et al. (Salem et al., 2022) used fine-tuning strategies based on a pre-trained model and achieved good classification results, which also reduced training samples by half. Zhao et al. (Zhao et al., 2020) proposed the double transfer method to solve the problem of sample scarcity. However, the long-tail distribution of object categories, where a few categories dominate and the majority are underrepresented, still poses a critical hurdle.
@@
Figure 1. Examples and data statistics of AIMO and RMO. (a): Examples of AIMO with multiple weather and illumination conditions. (b): Examples of RMO with multiple weather and illumination conditions. (c): The number of AIMO and RMO with multiple weather and illumination conditions. (d): The number of AIMO with different categories. (e): The number of RMO with different categories.
@@
The performance of maritime object recognition is also complicated by adverse weather conditions, such as fog, rain, and low visibility. DTDNet (Liu and Zhou, 2022) proposed an effective dehazing method to guarantee reliable ship detection under foggy conditions. IRDCLNet (Sun et al., 2022b) aimed to improve the ship instance segmentation performance on proposed Foggy ShipInsseg dataset. However, the methods trained on specific weather phenomena usually fail to generalize across various maritime conditions. FREGNet (Tian et al., 2024) concentrated on reducing the impact of adverse weather and uneven illumination on supervised object recognition. Raza et al. (Raza et al., 2022) designed SimuShips dataset consisting of maritime objects under multiple weather and time conditions, but this work lacks specific categories for objects, which is only used for obstacle object detection. In summary, unsupervised maritime object classification under various weather conditions has not been well addressed yet, and new approaches need to be explored.
@@
Due to the shortage of a large label-rich and balanced maritime object data under various weather and illumination conditions, we create AI-generated dataset for Maritime Object classification (AIMO), to compensate for the scarcity of weather conditions in real datasets and address the imbalance in the number of maritime objects across different categories. As shown in Fig. 1, AIMO is generated by Stable Diffusion, and naturally has a rich set of labels for categories and labels for multiple weather and illumination conditions. We also construct a real image dataset RMO (Real-world Maritime Object benchmark) by collecting from existing datasets and website. Even though advanced diffusion models generate images that are visually indistinguishable from real images, there are still significant differences between their distributions (You et al., 2024). Therfore, we propose a domain adaptation framework in which the generated data serves as the source domain, and the unlabeled real-world data acts as the target domain, facilitating improved classification accuracy under diverse weather conditions in real-world scenarios.
@@
Considering that the generated data may suffer from insufficient diversity and poor generalization, we first enhance the feature representation of the source domain using CLIP (Radford et al., 2021). Specifically, we design text prompts by combining maritime object categories and weather conditions, and train an image feature extractor to be aligned with the image and text feature spaces of CLIP. Facing the unreliability of the generated source domain and the unlabeled target domain, inspired by Sun et al. (Sun et al., 2022a), we introduce adaptive perturbations to the features to prevent overfitting during training. Furthermore, we employ a dynamic curriculum learning strategy. We comprehensively consider AIMO’s weather and category information, and gradually select samples for training to make model easy to converge.
@@
2. Related Work
2.1. AI-generated Datasets
@@
Researcher at MIT discovered that, for the person capture task, models trained on generated data (videos with fewer background objects) outperform those trained on real data (Zewe, 2022). He et al. (He et al., 2023) identified that larger data volumes and greater data diversity are crucial for achieving better pre-training outcomes with generated data. Furthermore, from the perspectives of model architecture and pre-training methodologies, Vision Transformer (ViT)-based models tend to be more effective for pre-training on generated data than convolutional neural networks (CNNs). Additionally, self-supervised learning methods have been shown to be more suitable for pre-training with generated data than traditional supervised methods (He et al., 2023).
@@
However, recent studies indicate that the use of AI-generated data may lead to model collapse (Shumailov et al., 2024). It is probably due to the generated data not being able to adequately represent the diversity and distribution of real data, which causes the model to gradually overfit these distorted data during training, ultimately compromising its ability to generalize, and leading to model collapse in the worst case. Model collapse not only diminishes accuracy, but also fosters self-reinforcing biases within the model, which further undermines the reliability of models. In light of this challenge, our research aims to utilize AI-generated data gradually from easy to difficult samples. We leverage the generated dataset, AIMO, for an unsupervised framework based on ViT in maritime object classification. Also, we utilized an Adaptive Adjustment Mechanism to address the issues of data overfitting. To the best of our knowledge, we are the first to overcome the long-tail issue in aspects of object categories and weather conditions for unsupervised maritime object classification by leveraging AI-generated data.
2.2. Unsupervised Domain Adaptation
@@
Unsupervised domain adaptation (UDA) seeks to align the source and target domains by learning a feature representation that is invariant across domains. A series of method for align domains is to minimize the distribution divergence between source and target domains with discrepancy measures, i.e., maximum mean discrepancy (MMD) (Kang et al., 2019), central moment discrepancy (CMD) (Zellinger et al., 2017), correlation alignment (CORAL) (Sun et al., 2016) and maximum density divergence (MDD) (Li et al., 2020). Another kind of methods is motivated by adversarial learning, playing a two-player min-max game to learn domain-invariant representations, i.e. DANN (Ganin and Lempitsky, 2015), which introduced a domain discriminator to generate domain-invariant features. Moreover, self-training is used for domain adaptation to generate pseudo-label for the target domain and apply it in the training process (Mei et al., 2020).
@@
With the rise of powerful Vision Transformer (ViT) for global feature extraction and modeling, an increasing number of research efforts are focusing on the application of transformers in UDA. ViT-base and ViT-small are from a pioneering work (Dosovitskiy et al., 2021) applying ViT to the task of image classification. Many works have been proposed on this basis. Swin (Liu et al., 2021), (Liu et al., 2022) focused on local self-attention computation and further proposed a large-scale pre-trained model serving as a generic backbone for computer vision. PMTrans (Zhu et al., 2023) used an intermediate domain to connect the source domain and the target domain, in order to bridge the domain gap, and conducted a min-max cross-entropy game on it. CDTrans (Xu et al., 2021) proposed cross–attention on source-target image pairs for direct feature alignment. TVT (Yang et al., 2023) proposed a multi-head self-attention module to obtain both transferable and discriminative features and combined it with adversarial adaptation.
@@
Moreover, considering the strong generalization ability of Vision-Language Models (VLMs), a series of methods focus on the combination of UDA with VLMs. It is proved that VLMs are capable of handling multi-modal information (text and image) to further minimizes the domain gap. Many works have focused on fine-tuning the VLMs. Cheng et al. (Cheng et al., 2024) subsequently disentangled text prompts into domain-invariant and domain-specific descriptions. Li et al. (Li et al., 2024) designed a modality separation network to distinctly disentangle VLM’s image embeddings into language-associated and vision-associated components. However, directly fine-tuning these VLMs into downstream tasks may be computationally expensive, due to the large number of parameters (Lai et al., 2024). Prompt tuning is an efficient learning method with freezing both image and text encoders. Lai et al. (Lai et al., 2024) proposed a domain-aware pseudo-labeling scheme designed for VLMs to achieve domain disentanglement. Bai et al. (Bai et al., 2024) proposed a Prompt-based Distribution Alignment (PDA) method to integrate the domain knowledge into prompt learning. DAMP (Du et al., 2024) learned prompts mutually with a cross-attention module based on text embeddings and image embeddings extracted from VLM.
@@
Figure 2. An overview of the proposed method. We use a series of Vision Transformer Blocks as backbone of feature extraction and take labels (classes and weather conditions) and images from the source domain, as well as images from the target domain as inputs. The designed UDA framework for maritime object classification is an adversarial adaptation network, consisting of generalization enhancement for source features, adversarial adaptation from source to target features and classifier refinement with perturbed features from both domains.
3. Method
@@
Specifically, the Generalization Enhancement aligns the features from Transformer Blocks with the text-encoder and image-encoder features of CLIP for source domain, which could be regarded as a Self-Knowledge Distillation process. The inputs of CLIP are the source domain data with rich information of category and multiple weather conditions. We measure the similarity between text embeddings, image embeddings and source domain features to constrain this process. Inspired by Sun et al. (Sun et al., 2022a), Classifier refinement improves the prediction accuracy by enhancing the robustness of feature representation. Random offsets are added into the input token sequence of a randomly chosen transformer block, which is performed on images from both source and target domains to acquire perturbed features. Then we employ the Kullback-Leibler (KL) divergence to quantify the discrepancy of the corresponding predicted class probabilities of original and perturbed features, and carry out random offset refinement. Furthermore, we use an Adaptive Adjustment Mechanism to carry out the training process step by step by scoring the source domain data.
3.1. Preliminary
@@
The basic adversarial UDA framework is composed of three parts: feature extractor F, consisting of a Patch Embedding layer and a series of Transformer Blocks, classifier C, and discriminator D. Given an input $\bm{x}$ from either source domain or target domain, $\bm{S}=\textbf{{F}}(\bm{x})$ is referred to a token sequence encoded by F, which is passed to D directly. Our novel adversarial UDA framework contains a Generalization Enhancement module G for enhancing the source domain feature. Therefore, the source domain feature $\bm{S}_{\bm{x}^{s}}$ is passed to G, and then $\bm{S}^{\textbf{{G}}}_{\bm{x}^{s}}=\textbf{{G}}(\bm{S}_{\bm{x}^{s}})$ is passed to C. The C predict $\bm{S}^{\textbf{{G}}}_{\bm{x}^{s}}$ into class probabilities $\bm{p}=\textbf{{C}}(\bm{S}^{\textbf{{G}}}_{\bm{x}^{s}})\in\mathcal{R}^{K}$. Therefore, the classification loss on source domain data is computed as
(1) $\mathcal{L}_{cls}=-\mathbb{E}_{(\bm{x},\bm{y})\in\mathcal{D}_{s}}\sum_{k=1}^{K}\mathbb{1}_{k=y}\log{\bm{p}_{k}},$
which is actually a standard cross-entropy loss.
@@
(2) $\mathcal{L}_{dom}=-\mathbb{E}_{\bm{x}\in\mathcal{D}_{s}}\log{\bm{q}_{s}}-\mathbb{E}_{\bm{x}\in\mathcal{D}_{t}}\log{\bm{q}_{t}},$
@@
(3) $\min_{\textbf{{F}},\textbf{{C}}}\max_{\textbf{{D}}}\mathcal{L}=\mathcal{L}_{cls}-\mathcal{L}_{dom}+\alpha\mathcal{L}_{SKD}+\beta\mathcal{L}_{offset},$
@@
where $\alpha$ and $\beta$ are trade-off parameters introduced in Sec 3.4. $\mathcal{L}_{SKD}$ represents the Self-Knowledge Distillation (SKD) loss introduced in Sec 3.2 and $\mathcal{L}_{offset}$ represents the random offset refinement loss on both source and target domain introduced in Sec 3.3.
3.2. Generalization Enhancement Based on CLIP

### v3

@@
Abstract
The classification and recognition of maritime objects are crucial for enhancing maritime safety, monitoring, and intelligent sea environment prediction. However, existing unsupervised methods for maritime object classification often struggle with the long-tail data distributions in both object categories and weather conditions. In this paper, we construct a dataset named AIMO produced by large-scale generative models with diverse weather conditions and balanced object categories, and collect a dataset named RMO with real-world images where long-tail issue exists. We propose a novel unsupervised domain adaptation approach that leverages AIMO (source domain) to address the problem of limited labeled data, unbalanced distribution and domain shift in RMO (target domain). Specifically,we enhance the generalization of source features with the Vision-Language Models (i.e., CLIP) by Self-Knowledge Distillation. And we propose a difficulty score for curriculum learning to optimize training process by gradually introducing the generated data. Experimental results shows that the proposed method significantly improves the classification accuracy, particularly for samples within rare object categories and weather conditions. Datasets and codes will be publicly available at https://github.com/honoria0204/AIMO.
keywords:
[…]
1 Introduction
@@
Unsupervised maritime object classification presents significant challenges in handling the high similarity between ships and complex maritime environments. Large-scale annotated datasets are not common because of the difficulty in obtaining maritime objects and the high cost of manual labeling wang2020generalizing. Salem et al. salem2022transfer used fine-tuning strategies based on a pre-trained model and achieved good classification results, which also reduced training samples by half. Zhao et al. zhao2020maritime proposed the double transfer method to solve the problem of sample scarcity. Zhang et al. zhang2025moment matched the joint distributions across domains, independent of the availability of source data. Zhang et al. DBLP:journals/ipm/ZhangZHM24 considered that the target data are usually highly heterogeneous and designed a Cross-domain Knowledge Collaboration to deal with mixed multiple target domains. However, the long-tail distribution of object categories, where a few categories dominate and the majority are underrepresented, still poses a critical hurdle.
@@
Figure 1: Examples and data statistics of AIMO and RMO. (a): Examples of AIMO with multiple weather and illumination conditions. (b): Examples of RMO with multiple weather and illumination conditions. (c): The number of AIMO and RMO with multiple weather and illumination conditions. (d): The number of AIMO with different categories. (e): The number of RMO with different categories.
@@
The performance of maritime object recognition is also complicated by adverse weather conditions, such as fog, rain, and low visibility. DTDNet liu2022dual proposed an effective dehazing method to guarantee reliable ship detection under foggy conditions. IRDCLNet sun2022irdclnet aimed to improve the ship instance segmentation performance on proposed Foggy ShipInsseg dataset. However, the methods trained on specific weather phenomena usually fail to generalize across various maritime conditions. FREGNet tian2024fregnet concentrated on reducing the impact of adverse weather and uneven illumination on supervised object recognition. Raza et al. raza2022simuships designed SimuShips dataset consisting of maritime objects under multiple weather and time conditions, but this work lacks specific categories for objects, which is only used for obstacle object detection. In summary, unsupervised maritime object classification under various weather conditions has not been well addressed yet, and new approaches need to be explored.
@@
Due to the shortage of a large label-rich and balanced maritime object data under various weather and illumination conditions, we create AI-generated dataset for Maritime Object classification (AIMO), to compensate for the scarcity of weather conditions in real datasets and address the imbalance in the number of maritime objects across different categories. As shown in Fig. 1, AIMO is generated by Stable Diffusion, and naturally has a rich set of labels for categories and labels for multiple weather and illumination conditions. We also construct a real image dataset RMO (Real-world Maritime Object benchmark) by collecting from existing datasets and website. Even though advanced diffusion models generate images that are visually indistinguishable from real images, there are still significant differences between their distributions you2024are. Therfore, we propose a domain adaptation framework in which the generated data serves as the source domain, and the unlabeled real-world data acts as the target domain, facilitating improved classification accuracy under diverse weather conditions in real-world scenarios.
@@
Considering that the generated data may suffer from insufficient diversity and poor generalization, we first enhance the feature representation of the source domain using CLIP radford2021learning. Specifically, we design text prompts by combining maritime object categories and weather conditions, and train an image feature extractor to be aligned with the image and text feature spaces of CLIP. Facing the unreliability of the generated source domain and the unlabeled target domain, inspired by Sun et al. sun2022safe, we introduce adaptive perturbations to the features to prevent overfitting during training. Furthermore, we employ a dynamic curriculum learning strategy. We comprehensively consider AIMO’s weather and category information, and gradually select samples for training to make model easy to converge.
@@
2 Related Work
2.1 AI-generated Datasets
@@
Researcher at MIT discovered that, for the person capture task, models trained on generated data (videos with fewer background objects) outperform those trained on real data zewe2022machine. He et al. DBLP:conf/iclr/HeS0XZTBQ23 identified that larger data volumes and greater data diversity are crucial for achieving better pre-training outcomes with generated data. Furthermore, from the perspectives of model architecture and pre-training methodologies, Vision Transformer (ViT)-based models tend to be more effective for pre-training on generated data than convolutional neural networks (CNNs). Additionally, self-supervised learning methods have been shown to be more suitable for pre-training with generated data than traditional supervised methods DBLP:conf/iclr/HeS0XZTBQ23.
@@
However, recent studies indicate that the use of AI-generated data may lead to model collapse shumailov2024ai. It is probably due to the generated data not being able to adequately represent the diversity and distribution of real data, which causes the model to gradually overfit these distorted data during training, ultimately compromising its ability to generalize, and leading to model collapse in the worst case. Model collapse not only diminishes accuracy, but also fosters self-reinforcing biases within the model, which further undermines the reliability of models. In light of this challenge, our research aims to utilize AI-generated data gradually from easy to difficult samples. We leverage the generated dataset, AIMO, for an unsupervised framework based on ViT in maritime object classification. Also, we utilized an Adaptive Adjustment Mechanism to address the issues of data overfitting. To the best of our knowledge, we are the first to overcome the long-tail issue in aspects of object categories and weather conditions for unsupervised maritime object classification by leveraging AI-generated data.
2.2 Unsupervised Domain Adaptation
@@
Unsupervised domain adaptation (UDA) seeks to align the source and target domains by learning a feature representation that is invariant across domains. A series of method for align domains is to minimize the distribution divergence between source and target domains with discrepancy measures, i.e., maximum mean discrepancy (MMD) kang2019contrastive, central moment discrepancy (CMD) zellinger2017central, correlation alignment (CORAL) sun2016return and maximum density divergence (MDD) li2020maximum. Another kind of methods is motivated by adversarial learning, playing a two-player min-max game to learn domain-invariant representations, i.e. DANN ganin2015unsupervised, which introduced a domain discriminator to generate domain-invariant features. Moreover, self-training is used for domain adaptation to generate pseudo-label for the target domain and apply it in the training process mei2020instance.
@@
With the rise of powerful Vision Transformer (ViT) for global feature extraction and modeling, an increasing number of research efforts are focusing on the application of transformers in UDA. ViT-base and ViT-small are from a pioneering work DBLP:conf/iclr/DosovitskiyB0WZ21 applying ViT to the task of image classification. Many works have been proposed on this basis. Swin liu2021swin, liu2022swin focused on local self-attention computation and further proposed a large-scale pre-trained model serving as a generic backbone for computer vision. PMTrans zhu2023patch used an intermediate domain to connect the source domain and the target domain, in order to bridge the domain gap, and conducted a min-max cross-entropy game on it. CDTrans xu2021cdtrans proposed cross–attention on source-target image pairs for direct feature alignment. TVT yang2023tvt proposed a multi-head self-attention module to obtain both transferable and discriminative features and combined it with adversarial adaptation.
@@
Moreover, considering the strong generalization ability of Vision-Language Models (VLMs), a series of methods focus on the combination of UDA with VLMs. It is proved that VLMs are capable of handling multi-modal information (text and image) to further minimizes the domain gap. Many works have focused on fine-tuning the VLMs. Cheng et al. cheng2024disentangled subsequently disentangled text prompts into domain-invariant and domain-specific descriptions. Li et al. li2024split designed a modality separation network to distinctly disentangle VLM’s image embeddings into language-associated and vision-associated components. However, directly fine-tuning these VLMs into downstream tasks may be computationally expensive, due to the large number of parameters lai2024empowering. Prompt tuning is an efficient learning method with freezing both image and text encoders. Lai et al. lai2024empowering proposed a domain-aware pseudo-labeling scheme designed for VLMs to achieve domain disentanglement. Bai et al. bai2024prompt proposed a Prompt-based Distribution Alignment (PDA) method to integrate the domain knowledge into prompt learning. DAMP du2024domain learned prompts mutually with a cross-attention module based on text embeddings and image embeddings extracted from VLM. DTKI zhan2026dual integrated local feature structures from CLIP with inter-class relationships from the source model to guide adaptation.
@@
Figure 2: An overview of the proposed method. We use a series of Vision Transformer Blocks as backbone of feature extraction and take labels (classes and weather conditions) and images from the source domain, as well as images from the target domain as inputs. The designed UDA framework for maritime object classification is an adversarial adaptation network, consisting of generalization enhancement for source features, adversarial adaptation from source to target features and classifier refinement with perturbed features from both domains.
3 Methodology
@@
Specifically, the Generalization Enhancement aligns the features from Transformer Blocks with the text-encoder and image-encoder features of CLIP for source domain, which could be regarded as a Self-Knowledge Distillation process. The inputs of CLIP are the source domain data with rich information of category and multiple weather conditions. We measure the similarity between text embeddings, image embeddings and source domain features to constrain this process. Inspired by Sun et al. sun2022safe, Classifier refinement improves the prediction accuracy by enhancing the robustness of feature representation. Random offsets are added into the input token sequence of a randomly chosen transformer block, which is performed on images from both source and target domains to acquire perturbed features. Then we employ the Kullback-Leibler (KL) divergence to quantify the discrepancy of the corresponding predicted class probabilities of original and perturbed features, and carry out random offset refinement. Furthermore, we use an Adaptive Adjustment Mechanism to carry out the training process step by step by scoring the source domain data.
3.1 Preliminary
@@
The basic adversarial UDA framework is composed of three parts: feature extractor F, consisting of a Patch Embedding layer and a series of Transformer Blocks, classifier C, and discriminator D. Given an input $\bm{x}$ from either source domain or target domain, $\bm{S}=\textbf{{F}}(\bm{x})$ is referred to a token sequence encoded by F, which is passed to D directly. Our novel adversarial UDA framework contains a Generalization Enhancement module G for enhancing the source domain feature. Therefore, the source domain feature $\bm{S}_{\bm{x}^{s}}$ is passed to G, and then $\bm{S}^{\textbf{{G}}}_{\bm{x}^{s}}=\textbf{{G}}(\bm{S}_{\bm{x}^{s}})$ is passed to C. The C predict $\bm{S}^{\textbf{{G}}}_{\bm{x}^{s}}$ into class probabilities $\bm{p}=\textbf{{C}}(\bm{S}^{\textbf{{G}}}_{\bm{x}^{s}})\in\mathcal{R}^{K}$. Therefore, we calculate the standard cross-entropy loss on source domain data as classification loss.
$\mathcal{L}_{ce}=-\mathbb{E}_{(\bm{x},\bm{y})\in\mathcal{D}_{s}}\sum_{k=1}^{K}\mathbb{1}_{k=y}\log{\bm{p}_{k}}.$ (1)
Considering the unbalanced samples, we further design a focal loss based on the cross-entropy loss.
$\mathcal{L}_{focal}=-\mathbb{E}_{(\bm{x},\bm{y})\in\mathcal{D}_{s}}\sum_{k=1}^{K}\mathbb{1}_{k=y}(1-\bm{p}_{k})^{\tau}\log{\bm{p}_{k}},$ (2)
when $\tau=0$, focal loss degenerates to cross-entropy loss. And the larger t is, the greater the weight of the inaccurate samples in the loss. We will detail the choice of the value of $\tau=0$ in Sec 3.4.
@@
$\mathcal{L}_{dom}=-\mathbb{E}_{\bm{x}\in\mathcal{D}_{s}}\log{\bm{q}_{s}}-\mathbb{E}_{\bm{x}\in\mathcal{D}_{t}}\log{\bm{q}_{t}},$ (3)
@@
$\min_{\textbf{{F}},\textbf{{C}}}\max_{\textbf{{D}}}\mathcal{L}=\mathcal{L}_{focal}-\mathcal{L}_{dom}+\alpha\mathcal{L}_{SKD}+\beta\mathcal{L}_{offset},$ (4)
@@
where $\alpha$ and $\beta$ are trade-off parameters introduced in Sec 3.4. $\mathcal{L}_{SKD}$ represents the Self-Knowledge Distillation (SKD) loss introduced in Sec 3.2 and $\mathcal{L}_{offset}$ represents the random offset refinement loss on both source and target domain introduced in Sec $3.$
3.2 Generalization Enhancement Based on CLIP